# Primacy Effect of ChatGPT

### v1

@@ Abstract
Instruction-tuned large language models (LLMs), such as ChatGPT, have led to promising zero-shot performance in discriminative natural language understanding (NLU) tasks. This involves querying the LLM using a prompt containing the question, and the candidate labels to choose from. The question-answering capabilities of ChatGPT arise from its pre-training on large amounts of human-written text, as well as its subsequent fine-tuning on human preferences, which motivates us to ask: Does ChatGPT also inherits humans’ cognitive biases? In this paper, we study the primacy effect of ChatGPT: the tendency of selecting the labels at earlier positions as the answer. We have two main findings: i) ChatGPT’s decision is sensitive to the order of labels in the prompt; ii) ChatGPT has a clearly higher chance to select the labels at earlier positions as the answer. We hope that our experiments and analyses provide additional insights into building more reliable ChatGPT-based solutions. We release the source code at https://github.com/wangywUST/PrimacyEffectGPT.
@@ 2.1 Prompts for ChatGPT
Prompts are a key component to the effective use of ChatGPT on discriminative NLU tasks Wei et al. (2023); Yuan et al. (2023). Generally, prompts for such tasks involves two key components: (i) label definitions, and (ii) a task description and input text (see an example in Fig. 2).
ChatGPT’s capability of understanding instructions in the prompt benefits from its training with human feedback OpenAI (2022), but this also creates the risks of inheriting humans’ cognitive biases. In this paper, we discuss a cognitive bias in ChatGPT: the primacy effect, which indicates the tendency of selecting labels in earlier positions in the prompt.
@@ 2.2 Analysis with Label Shuffling
{tcolorbox}
[fonttitle=,title=Prompts for Zero-shot Intent Detection, size = normal, label=mybox] Label Definitions
@@
Label 2: card arrival
Label 3: activate my card
[…]
Which Label matches the intent expressed in the Target Text?
@@ 4 Related Work
Large Language Models (LLMs) Brown et al. (2020); Rae et al. (2021); Thoppilan et al. (2022); Smith et al. (2022), such as GPT-3 Brown et al. (2020), LaMDA Thoppilan et al. (2022) and PaLM Chowdhery et al. (2022), refer to large scale pretrained model that contain more than a hundred billion parameters. Based on the the highly parallelizable Transformer architecture Vaswani et al. (2017), these Large Language models have shown powerful capability to produce reasonable results with very few samples or task descriptions as input.
A key milestone in the development process of LLMs is ChatGPT, which is developed by OpenAI based on InstructGPTOuyang et al. (2022). ChatGPT is able to interact with humans through multiple turns of dialogue, understand user intent, accomplish instructions and return human-like responses. This attracts huge attention from research field, motivating numerous recent work Zhang et al. (2022); Ma et al. (2023); Wan et al. (2023); Zhong et al. (2023); Susnjak (2023) to utilize ChatGPT to different tasks.
As ChatGPT is a proprietary model, and OpenAI does not disclose its training specifics, researchers are actively investigating its associated implications and capabilities. There have been some work analyzing the performance, robustness, faithfulness, and explain-ability of ChatGPT Gao et al. (2023); Han et al. (2023); Li et al. (2023). For example, Malinka et al. (2023) investigates the educational integrity of ChatGPT and evaluates the ChatGPT’s abilities to solve assignments of various levels in computer security specialization. Haque et al. (2022) and Krügel et al. (2023) investigate the ethical risks of ChatGPT.
@@
Different from the above work, we focus on a new phenomenon of ChatGPT: the primacy effect, which is the tendency of selecting the first labels as the answer. The Primacy effect seriously influence ChatGPT’s fairness. Collectively, our findings provide a new understanding of how ChatGPT works given the instructional prompts.
@@ 5 Conclusion
While previous work often takes ChatGPT as a universal method applicable to all text-related tasks, we argue that its flexibility comes with the risk of inheriting human’s cognitive biases. These biases lead to unfair judgments which can affect the performance of the machine learning model. This work studies a cognitive bias of ChatGPT: primacy effects. We propose a simple yet effective label shuffling method to analyze the influence of label orders on ChatGPT. We discover the primacy effects of ChatGPT and finds that it highly influence the fairness of ChatGPT in NLU. Our work contributes to better understanding the behaviors of ChatGPT and building more faithful solutions with ChatGPT in NLU applications.

### v2

@@ Abstract
Instruction-tuned large language models (LLMs), such as ChatGPT, have led to promising zero-shot performance in discriminative natural language understanding (NLU) tasks. This involves querying the LLM using a prompt containing the question, and the candidate labels to choose from. The question-answering capabilities of ChatGPT arise from its pre-training on large amounts of human-written text, as well as its subsequent fine-tuning on human preferences, which motivates us to ask: Does ChatGPT also inherit humans’ cognitive biases? In this paper, we study the primacy effect of ChatGPT: the tendency of selecting the labels at earlier positions as the answer. We have two main findings: i) ChatGPT’s decision is sensitive to the order of labels in the prompt; ii) ChatGPT has a clearly higher chance to select the labels at earlier positions as the answer. We hope that our experiments and analyses provide additional insights into building more reliable ChatGPT-based solutions. We release the source code at https://github.com/wangywUST/PrimacyEffectGPT.
@@ 2.1 Prompts for ChatGPT
Prompts are a key component to the effective use of ChatGPT on discriminative NLU tasks Wei et al. (2023); Yuan et al. (2023). Generally, prompts for such tasks involve two key components: (i) label definitions, and (ii) a task description and input text (see an example in Fig. 2).
ChatGPT’s capability of understanding instructions in the prompt benefits from its training with human feedback OpenAI (2022), but this also creates the risk of inheriting humans’ cognitive biases. In this paper, we discuss a cognitive bias in ChatGPT: the primacy effect, which indicates the tendency of selecting labels in earlier positions in the prompt.
@@ 2.2 Analysis with Label Shuffling
Prompts for Zero-shot Intent Detection Label Definitions
@@
Label 2: card arrival Label 3: activate my card
Target Text: I need a new PIN. Which Label matches the intent expressed in the Target Text?
@@ 4 Related Work
Large Language Models (LLMs) Brown et al. (2020); Rae et al. (2021); Thoppilan et al. (2022); Smith et al. (2022), such as GPT-3 Brown et al. (2020), LaMDA Thoppilan et al. (2022) and PaLM Chowdhery et al. (2022), refer to large scale pretrained models that contain more than a hundred billion parameters. Based on the highly parallelizable Transformer architecture Vaswani et al. (2017), these Large Language models have shown powerful capability to produce reasonable results with very few samples or task descriptions as input.
A key milestone in the development process of LLMs is ChatGPT, which is developed by OpenAI based on InstructGPTOuyang et al. (2022). ChatGPT is able to interact with humans through multiple turns of dialogue, understand user intent, accomplish instructions, and return human-like responses. This attracts huge attention from research field, motivating numerous recent work Zhang et al. (2022); Ma et al. (2023); Wan et al. (2023); Zhong et al. (2023); Susnjak (2023) to utilize ChatGPT to different tasks.
As ChatGPT is a proprietary model, and OpenAI does not disclose its training specifics, researchers are actively investigating its associated implications and capabilities. There has been some work analyzing the performance, robustness, faithfulness, and explain-ability of ChatGPT Gao et al. (2023); Han et al. (2023); Li et al. (2023). For example, Malinka et al. (2023) investigates the educational integrity of ChatGPT and evaluates the ChatGPT’s abilities to solve assignments of various levels in computer security specialization. Haque et al. (2022) and Krügel et al. (2023) investigate the ethical risks of ChatGPT.
@@
Different from the above work, we focus on a new phenomenon of ChatGPT: the primacy effect, which is the tendency of selecting the first labels as the answer. The primary effect seriously influences ChatGPT’s fairness. Collectively, our findings provide a new understanding of how ChatGPT works given the instructional prompts.
@@ 5 Conclusion
While previous work often takes ChatGPT as a universal method applicable to all text-related tasks, we argue that its flexibility comes with the risk of inheriting human’s cognitive biases. These biases lead to unfair judgments which can affect the performance of the machine learning model. This work studies a cognitive bias of ChatGPT: primacy effects. We propose a simple yet effective label shuffling method to analyze the influence of label orders on ChatGPT. We discover the primacy effect of ChatGPT and finds that it highly influences the fairness of ChatGPT in NLU. Our work contributes to a better understanding of the behaviors of ChatGPT and building more faithful solutions with ChatGPT in NLU applications.